1). Sales is $100 per unit; direct material is $22 per unit; direct labor is $18 per unit; overhead is $8 per unit and $1,800 per month; selling and administrative expenses are $5 per unit and $1,400 per month. If the company produced 100 units, total period costs are:

a).$1,900 b).$4,800 c).$6,600 d).$500

2).Sales is $100 per unit; direct material is $25 per unit; direct labor is $15 per unit, overhead is $10 per unit and $20,000 per month; selling and administrative expenses are $5 per unit and $10,000 per month. The company produced 900 units.  Contribution margin equals:

a).$40,500 b).$25,000 c).$10,500 d).$49,500

3).Sales is $120 per unit; direct material is $21 per unit; direct labor is $9 per unit, overhead is $6 per unit and $9,000 per month; selling and administrative expenses are $12 per unit and $18,000 per month. The company produced 1,000 units. Total product costs (cost of goods sold) are:

a).$45,000 b).$36,000 c).$30,000 d).$75,000
Answer #1

1) Total period Cost = (100*5+1400) = 1900

So answer is a) $1900

2) Contribution margin = Sales-Variable cost = (100-25-15-10-5)*900 = 40500

So answer is a) $40500

3) Total product cost = (21+9+6)*1000+9000 = 45000

So answer is a) $45000